# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v1

@@ 3 Generalized Cross-Validation
Before stating our results, we note the running assumptions used in what follows. We consider a sequence of right-rotationally invariant designs $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n},\dots$ where each $\mathbf{X}_{i}\sim\mathbf{Q}_{i}^{\top}\mathbf{D}_{i}\mathbf{O}_{i}$, and a sequence of signal vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\dots,\boldsymbol{\beta}_{n},\dots$ such that for all $n$, $\boldsymbol{\beta}_{n}\in\mathbb{R}^{p}$ and $\|\boldsymbol{\beta}_{n}\|/\sqrt{n}=r$. We generate $\mathbf{y}_{i}$ according to (1). In the sequel, we drop the dependence on $n$, since it is clear from context, and just write $\mathbf{X},\mathbf{y}$ and $\boldsymbol{\beta}$ with the indexing implicit. Furthermore, the following hold:
  1. A1
    lim $\limsup\lambda_{\max}(\mathbf{D}_{n})<C$ almost surely for some constant $C$.
  2. A2
    Each $\mathbf{X}_{n}$ has dimensions $n\times p(n)$, with $p(n)/n\xrightarrow{n\to\infty}\gamma\in[0,\infty)$.
  3. A3
    Each entry of $\boldsymbol{\epsilon}$ is independent with finite moment of order $4+\eta$, for some $\eta>0$. Furthermore, $\mathbb{E}[\epsilon_{i}]=0$ and $\mathbb{E}[\epsilon_{i}^{2}]=\sigma^{2}$.
  4. A4
    To fix scaling and simplify our results, we assume without loss of generality that $\frac{n^{\prime}\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{np}=1$.
  5. A5
    Each $\boldsymbol{\beta}_{n}$ is fixed or random independent of $\mathbf{X}_{n}$.
  6. A6
    $\widetilde{\mathbf{X}}$ is an independent draw from a potentially different right-rotationally invariant ensemble.
In all that follows, we state only the main results and defer proofs to the appendix.
3.1 An Analysis of Cross-Validation Methods
In [31], it is shown under a broad class of designs with i.i.d. rows that both LOOCV and GCV are asymptotically (under proportional asymptotics) consistent for the out-of-sample error. In our case, LOOCV is not tractable to analyze without additional assumptions on the distribution of $\mathbf{Q}$ – see Appendix B.2 for details. However, analysis of GCV, which was proposed as a rotationally invariant alternative to the standard leave-one-out cross-validation, is tractable. In particular, the GCV cross-validation metric is as follows:
$\mathsf{GCV}_{n}(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}}{1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n}\right)^{2}=\frac{1}{n-\mathsf{Tr}(\mathbf{S}_{\lambda})}\sum_{i=1}^{n}\left(y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}\right)^{2}.$ (4)
Here $\mathbf{S}_{\lambda}$ is the smoother matrix given by $\mathbf{X}^{\top}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}$. Our first result is as follows:
@@ 3.1 An Analysis of Cross-Validation Methods
We also find that for right-rotationally invariant designs, the asymptotic risk takes the following form:
@@
Fix $0<\lambda_{1}<\lambda_{2}<\infty$. Under the stated assumptions,
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}\left|R_{\mathbf{X},\mathbf{y}}\left(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta}\right)-\underbrace{\left(r^{2}\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)+\sigma^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))\right)}_{\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})}\right|\xrightarrow{a.s.}0.$ (6)
@@
The classical GCV then produces consistent tuning results if asymptotic value of GCV as a function of $\lambda$ (cf. Theorem 1) equals to the asymptotic risk as a function of $\lambda$ (cf. Theorem 2) up to some monotonic transform. This is what is observed for i.i.d. data under relatively lax assumptions, shown in [19] for the isotropic setting with random $\boldsymbol{\beta}$ and more generally in [31], but these proofs rely explicitly on identities given by the Silverstein equation [36, 19, 31]. Such identities do not exist outside of this i.i.d. setting, and the GCV is naturally biased once the data violates this assumption. Precisely, the asymptotic value of $\mathsf{GCV}_{n}$ in Theorem 1 is not necessarily equal to the asymptotic risk $\mathsf{R}_{\mathbf{D}}$ in Theorem 2 (up to an additive factor of $\sigma^{2}$) outside of the i.i.d. setting. We present an alternative approach which is provably consistent for right-rotationally invariant designs.
@@ 3.2 A Modified GCV
To obtain a consistent estimate of $R_{\mathbf{X},\mathbf{y}}$, our idea is to estimate $r^{2}$ and $\sigma^{2}$ and use these to compute the expression $\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})$ given by Theorem 2. We begin by discussing how estimate these two parameters.
@@ 3.2.1 Estimators for $r^{2},\sigma^{2}$
Estimating $\sigma^{2}$ in these settings has already been studied in [27], wherein a consistent estimator for $\sigma^{2}$ is presented. Our approach, which we observe to have better finite sample performance, will be slightly different. The original GCV can be seen as starting with the training error and finding a rescaling that recovers the out-of-sample risk. We likewise begin with the training error, but instead use it to produce estimating equations. In particular, we establish the following result:
@@
Under the stated assumptions, for any $\lambda>0$,
$\frac{1}{n}\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{y}-\left[r^{2}\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)+\sigma^{2}\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)\right]\xrightarrow{a.s.}0.$ (7)
Observe that equation (7) can be used as an estimating equation, meaning that, if one treats  (7) as an equality to 0 (i.e. replacing “$\xrightarrow{a.s.}$” with “$=$”), then we can solve for the the only unknown quantities are $r^{2}$ and $\sigma^{2}$, allowing us to solve for them. The first term, $\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{y}/n$ is the training error, then $\gamma=p/n$ along with $v_{\mathbf{D}}$ and $v^{\prime}_{\mathbf{D}}$, defined in Section 2.1, are all data dependent quantities. Moreover, it is an estimating equation for every value of $\lambda$, and each value yields a different estimating equation. While it takes only two values of $\lambda$ for us to solve for $r^{2}$ and $\sigma^{2}$, we find that this method is quite sensitive to the choice of $\lambda$. Instead, we choose a grid of $\lambda$ and find OLS estimates for $r^{2}$ and $\sigma^{2}$, which reduces this sensitivity. Explicitly, the scheme proceeds as shown at the start of Algorithm 1.
A simple corollary of Lemma 2 is then the following:
@@
$\hat{r}^{2}$ and $\hat{\sigma}^{2}$ are strongly consistent for $r^{2}$ and $\sigma^{2}$, respectively.
It should be possible to show that these estimators have $O(\sqrt{n})$ rates under stronger assumptions on the distribution of the noise, such as sub-Gaussianity, but this analysis is not pursued.
@@
1:$\mathbf{X},\mathbf{y}$, regularization strength $\lambda$; for use in estimation, grid of regularization strengths $\lambda_{1},\dots,\lambda_{L}$;
@@
9:return $\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$
@@ 3.2.2 Plug-in estimation
We now simply define $\mathsf{ROTI}\text{-}\mathsf{GCV}(\lambda)=\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$ (recall $\mathsf{R}_{\mathbf{D}}$ from Theorem 2) and tune $\lambda$ using this metric; furthermore, this quantity serves as an estimator for the out-of-sample risk. It is then clear from Theorem 2 that this is consistent on compact intervals.
@@
Fix $0<\lambda_{1}<\lambda_{2}<\infty$. Under Assumptions A1-A6, one has
@@
Replacing $\hat{r}^{2}$ and $\hat{\sigma}^{2}$ with any other set of consistent estimators also produces a valid estimate for the risk.
@@ 4 GCV under Signal-PC Alignment
An issue that emerges when attempting to apply GCV for right-rotationally invariant ensembles in practice is that the signal often tends to align with the top eigenvectors of the data, violating the independence assumption imposed on $\boldsymbol{\beta}$ and $\mathbf{O}$. The right-rotationally invariant assumption, for fixed $\boldsymbol{\beta}$, inherently assumes that the signal is incoherent with respect to the eigenbasis. However, this assumption is generally not true, and in fact in the i.i.d. anisotropic case, the geometry of $\boldsymbol{\beta}$ with respect to the top eigenvectors of $\boldsymbol{\Sigma}$ is known to significantly influence the behavior of ridge regression (see e.g. [45]).
Adapting the approach given in [27], we model this scenario using two index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ of distinguished eigenvectors, which are aligned and coupled eigenvectors, respectively. We replace the last two assumptions in Section 3 with the following:
  1. A7
    The true signal $\boldsymbol{\beta}$ is given as $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\sqrt{n}\alpha_{i}\mathbf{o}_{i}$, where $\|\boldsymbol{\beta}\|/\sqrt{n}=r$. An eigenvector $\mathbf{o}_{i}$ is an aligned eigenvector if $\mathbf{o}_{i}\in\mathcal{J}_{a}$, as then $\mathbf{o}_{i}$ aligns with the signal $\boldsymbol{\beta}$.
@@
    The test data $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$ is drawn from a separate right-rotationally invariant ensemble, where $\widetilde{\mathbf{O}}$ is distributed according to a Haar matrix where the eigenvectors are conditioned to satisfy $\mathbf{o}_{i}=\widetilde{\mathbf{o}}_{i}$ for all $i\in\mathcal{J}_{c}$. An eigenvector $\mathbf{o}_{i}$ is then coupled if $i\in\mathcal{J}_{c}$, as $\mathbf{o}_{i}$ is shared between the test and train sets.
@@
  1. A9
    $\mathcal{J}_{a},\mathcal{J}_{c}$, and the $\boldsymbol{\alpha}=(\alpha_{i})_{i\in\mathcal{J}_{a}}$ are fixed and do not change with $n,p$.
@@ 2. A10
    lim $\limsup\|\mathbb{E}[\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}]\|_{\mathsf{op}}<C$ for some constant $C$.
The main difficulty that arises in our setting, compared to [27], is that we must account for how the geometry of the eigenvectors of the test set relates to those of the training set. One should expect, for example, in some types of structured data that the top eigenvectors of the training data and test data are closely aligned. We account for this effect through the coupled eigenvectors condition (Assumption A8). Furthermore, note that without this coupling, any alignment between the eigenvectors of the training set and the signal $\boldsymbol{\beta}$ will not exist in the test set, since then the eigenvectors of the test set would be independent of this alignment. We are still interested in the same test risk $R_{\mathbf{X},\mathbf{y}}$ as in 2, except now note that $\widetilde{\mathbf{X}}$ and $\mathbf{X}$ are not independent.
@@
Under Assumptions A1-A4 and A7-A10, for any $0<\lambda_{1}<\lambda_{2}<\infty$, we have
@@
Denote $\mathfrak{d}_{i}^{2}=\mathbb{E}\left[\left(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}\right)_{ii}\right]$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}=\frac{1}{p-|\mathcal{J}_{c}|}\sum_{i\not\in\mathcal{J}_{c}}\mathfrak{d}_{i}^{2}$. Then
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}}$
$\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\sigma^{2}}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right]$
@@
The assumption that the top eigenvectors of training and test sample covariance are exactly equal is unrealistic (Assumption A8) and is not expected to hold in practice. However, the idea is that the approximation can lead to more robust procedures. The covariance structure imposed on the test data under the assumption of coupled eigenvectors takes the form $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$ (computed using [33, Lemma 4]). This can be interpreted as a low-rank approximation of the covariance of the test data with its bulk filled in with an identity piece. We thus expect this approximation to hold when the training and test data have “spiky” spectra, where the top few eigenvalues are heavily separated from the bulk, and these top eigenvectors of the test and training data are close.

### v2

@@ 3 Generalized Cross-Validation
Prior work [45, 29] established that for many designs with i.i.d. samples, LOOCV and GCV estimate the out-of-sample error consistently under the proportional asymptotics regime, while $k$-fold CV does not. In our setting, LOOCV is intractable without additional assumptions on $\mathbf{Q}$ (see discussion in Appendix B.6). However, GCV is tractable, motivating our study here. Initially introduced as a technique for selecting the regularization parameter in smoothing splines [6], GCV’s optimality has been shown in various tasks [21, 15]. It is also a consistent tuning technique for kernel ridge regression [47, 43, 44].
The GCV cross-validation metric for ridge regression is as follows [13]:
$\mathsf{GCV}_{n}(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}}{1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n}\right)^{2}.$ (4)
Here $\mathbf{S}_{\lambda}=\mathbf{X}^{\top}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}$ is the smoother matrix. Our first result is as follows:
@@
Further, the risk takes the following form:
@@
Fix $0<\lambda_{1}<\lambda_{2}<\infty$. Define
$\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})=r^{2}\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)+\sigma^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))$ (6)
Under the stated assumptions,
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}\left|R_{\mathbf{X},\mathbf{y}}\left(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta}\right)-\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})\right|\xrightarrow{a.s.}0.$ (7)
@@
In Theorem 2, one can interpret the term involving $r^{2}$ as the bias of the estimator, and the term with $\sigma^{2}$ as its variance. As such, one can then interpret Theorem 1 as showing how GCV reconstructs the bias and variance from the data. See Appendix A.4 for explicit derivations of the bias and variance terms.
Remark 3.
It is shown in [29] that for i.i.d. designs, $\Delta_{n}(\lambda):=\mathsf{GCV}_{n}(\lambda)-R_{\mathbf{X},\mathbf{y}}(\lambda)-\sigma^{2}$ approaches zero uniformly over compact intervals [^1] The extra $\sigma^{2}$ term is discussed in Remark 1.. Their proof of this relies explicitly on identities of the Stieltjes transform given by the Silverstein equation [33]. These identities do not exist for right-rotationally invariant designs, and thus the GCV becomes biased for these designs, meaning $\Delta_{n}(\lambda)\not\to 0$. The more the spectrum departs from that of the i.i.d. setting, the more biased the estimates.
@@ 3.1 A Modified GCV
The original GCV (in Eq. (4)) can be seen as starting with the training error and finding a rescaling that recovers the out-of-sample risk. To construct our alternative GCV metric, which is provably consistent for right-rotationally invariant designs, we likewise begin with the training error, but instead use it to produce estimating equations for $r^{2}$ and $\sigma^{2}$. We then use these to compute the risk formula given by Theorem 2.
@@ 3.1.1 Estimators for $r^{2},\sigma^{2}$
Consistent estimation of $\sigma^{2}$ has been studied in [23]. We propose a different approach since we observe it performs better in finite samples. In particular, we establish the following result:
@@
Under our assumptions, for any $\lambda>0$,
$\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}\right)^{2}-\left[r^{2}\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)+\sigma^{2}\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)\right]\xrightarrow{a.s.}0.$ (8)
Observe that Eq. (8) can be used as an estimating equation, meaning that, if one treats (8) as an equality to 0 (i.e. replacing “$\xrightarrow{a.s.}$” with “$=$”), then we can solve for the unknown quantities $r^{2}$ and $\sigma^{2}$.
Recall that $\gamma=p/n$, $v_{\mathbf{D}}$, and $v^{\prime}_{\mathbf{D}}$, defined in Section 2.1, are all data dependent quantities.
Now, Eq. (8) yields a distinct estimating equation for each value of $\lambda$. While it takes only two equations to solve for $r^{2}$ and $\sigma^{2}$, we instead use a grid of $\lambda$ and find OLS estimates for $r^{2}$ and $\sigma^{2}$ to improve robustness. Explicitly, the scheme proceeds as shown in steps 1-6 of Algorithm 1.
Using Lemma 1, we can establish that the following:
@@
Under Assumptions A1-A7 and a non-degeneracy condition on $\mathbf{D}$ (see Appendix A.7.1 for details), $\hat{r}^{2}$ and $\hat{\sigma}^{2}$ from Algorithm 1 are strongly consistent for $r^{2}$ and $\sigma^{2}$, respectively.
Empirically, we find that normalizing the spectrum of $\mathbf{X}^{\top}\mathbf{X}$ and taking $\{\lambda_{i}\}_{i=1}^{L}$ logarithmically spaced between $1$ and $10^{2.5}$ works well.
@@
1:$\mathbf{X},\mathbf{y}$, $\lambda$; for use in estimation, grid of regularization strengths $\lambda_{1}<\lambda_{2}\cdots<\lambda_{L}$;
@@
9:return $\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$ for $\mathsf{R}_{\mathbf{D}}(\cdot,\cdot)$ defined in (6)
@@ 3.1.2 Plug-in estimation
We now define $\mathsf{ROTI}\text{-}\mathsf{GCV}(\lambda):=\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$ (recall $\mathsf{R}_{\mathbf{D}}$ from Theorem 2) and tune $\lambda$ using this metric (Algorithm 1); below we establish that this quantity is uniformly consistent for the out-of-sample risk over compact intervals.
@@
Fix any $0<\lambda_{1}<\lambda_{2}<\infty$. Under Assumptions A1-A7, one has
@@
Uniform convergence, unlike pointwise convergence, ensures the risk attained by the minimizer (over a compact interval) of our cross-validation metric is close to optimal. See Appendix B.4 for a short proof.
@@ 4 GCV under Signal-PC Alignment
An issue that emerges when attempting to apply ROTI-GCV in practice is that the signal often tends to align with the top eigenvectors of the data, violating the independence assumption imposed on $\boldsymbol{\beta}$ and $\mathbf{O}$. The right-rotationally invariant assumption, for fixed $\boldsymbol{\beta}$, inherently assumes that the signal is incoherent with respect to the eigenbasis. However, this assumption is generally not true, and in fact in the i.i.d. anisotropic case, the geometry of $\boldsymbol{\beta}$ with respect to the top eigenvectors of $\boldsymbol{\Sigma}$ is known to significantly influence the behavior of ridge regression (see e.g. [42]).
Adapting the approach given in [23], we model this scenario using two index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ of distinguished eigenvectors, which are aligned and coupled eigenvectors, respectively. We replace Assumptions A6 and A7 in Section 3 with the following:
@@
    The true signal $\boldsymbol{\beta}$ is given as $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\sqrt{n}\alpha_{i}\mathbf{o}_{i}$, where $\|\boldsymbol{\beta}^{\prime}\|/\sqrt{n}=r$; $\boldsymbol{\beta}^{\prime}$ is independent of $(\mathbf{X},\boldsymbol{\epsilon})$. An eigenvector $\mathbf{o}_{i}$ is an aligned eigenvector if $\mathbf{o}_{i}\in\mathcal{J}_{a}$, as then $\mathbf{o}_{i}$ aligns with the signal $\boldsymbol{\beta}$;
  2. A9
    The test data $\widetilde{\mathbf{X}}\sim\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$ is drawn from a separate right-rotationally invariant ensemble; $\widetilde{\mathbf{O}}$ is distributed according to a Haar matrix with rows conditioned to satisfy $\mathbf{o}_{i}=\widetilde{\mathbf{o}}_{i}$ for all $i\in\mathcal{J}_{c}$. An eigenvector $\mathbf{o}_{i}$ is then coupled if $i\in\mathcal{J}_{c}$, as $\mathbf{o}_{i}$ is shared between the test and train sets.
@@ 1. A10
    $\mathcal{J}_{a},\mathcal{J}_{c}$, $\boldsymbol{\alpha}=(\alpha_{i})_{i\in\mathcal{J}_{a}}$ are fixed, not changing with $n,p$; this ensures consistent estimation of these distinguished directions.
  2. A11
    lim $\limsup\|\mathbb{E}[\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}]\|_{\mathsf{op}}<C$ for some constant $C$. This differs from Assumption A2, as it is on the test data.
The main difficulty that arises in our setting, compared to [23], is that we must account for how the geometry of the eigenvectors of the test set relates to those of the training set. One should expect, for example, in some types of structured data that the top eigenvectors of the training data and test data are closely aligned. We account for this effect through the coupled eigenvectors condition (Assumption A9). Furthermore, note that without this coupling, any alignment between the eigenvectors of the training set and the signal $\boldsymbol{\beta}$ will not exist in the test set, since then the eigenvectors of the test set would be independent of this alignment. We are still interested in the same test risk $R_{\mathbf{X},\mathbf{y}}$ as in Eq. (2), except now note that $\widetilde{\mathbf{X}}$ and $\mathbf{X}$ are not independent.
@@ 2. A11
Under Assumptions A1-A5 and A8-A11, for any $0<\lambda_{1}<\lambda_{2}<\infty$, we have
@@
Let $\mathfrak{d}_{i}^{2}=\mathbb{E}\left[\left(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}\right)_{ii}\right]$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}=\frac{1}{p-|\mathcal{J}_{c}|}\sum_{i\not\in\mathcal{J}_{c}}\mathfrak{d}_{i}^{2}$.
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\left[\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right]$
$\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\sigma^{2}}{n}\sum_{i=1}^{n}\frac{D_{ii}^{2}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}}$
@@
Comparing the two terms here with those in Theorem 1, observe that they are equal when $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ are empty. Furthermore, observe that each element of $\mathcal{J}_{a}$ adds a large contribution to the bias, scaling with the size of the alignment $\alpha_{i}$; similarly, elements of $\mathcal{J}_{c}$ contribute additional bias and variance.
Remark 6.
The assumption that the top eigenvectors of training and test sample covariance are exactly equal (Assumption A9) is not expected to hold perfectly in practice. However, the idea is that the approximation can lead to more robust procedures. We thus expect this approximation to hold when the training and test data have “spiky” spectra, where the top few eigenvalues are heavily separated from the bulk, and these top eigenvectors of the test and training data are close.